(x). (15 pts) to prove or disprove sqrt(2 + sqrt(6x)) = the definition of
Answer #1

Eriven: V2+ vox = √2+(6x)/2 (252+(6x)%]} = 2%2+ (446x14 Poredominant term in the above is 24 : V2tV6X = O(x4)